40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 388     Entry time: Fri Jan 27 17:22:34 2012
Author: Giordon Stark 
Type: DailyProgress 
Category: Coating Q 
Subject: Laser Issues (part 2) 

 The 1103P we found was pretty crappy and wasn't staying on. We broke for lunch, Zach had a meeting, and met up again around 3:30pm. Zach found another working one from Koji's office in Bridge and we set-up roughly what the experiment is and just checked out the equipment and the signals from the photodetectors, lock-in amplifier, pre-amps, and so on.

We ran a noise measurement on the photodiodes with and without the laser (they're the same, so it's just one measurement really, but we'll include both in a later ELOG post to show that they are the same). Quick observations (number crunching to be done in near future): we ran it from 0-100kHz and notice a rather broad (50kHz) hump around 60kHz which is most likely due to specs about the photodetectors themselves.

The noise of the pre-amplifier seems much higher than we expected - and it happens to be the cause of the bump - we grounded the input of the pre-amp and noticed the hump is there. We also increase the amplification by a factor of 10 and noticed that the noise floor increased to be at the same level as the hump - which tells us that the noise is introduced into the pre-amp after the amplification of the signal. We replaced the cable and noticed no change. It appears that by switching from power supply to battery power fixes this.

ELOG V3.1.3-